OpenStudy (anonymous):

kenny spent one hour more than lucas studying, andrea spent one hour less than twice the hours lucas spent. if together they spent 14 hours, how many more hours did lucas study?

OpenStudy (anonymous):

LET lucas = x, kenny = x+1, andrea = 2x-1. Then add them all together, and equate to 14.\[4x = 14\]\[x = 7/2\] or 3 1/2 hours

OpenStudy (anonymous):

translate to an equation do not solve

OpenStudy (anonymous):

K = L +1 (1), A = 2L -1 (2), K + L + A = 14 (3) From (1) and (3) L+1 = 14 - L - A => A = -2L + 13 (4) From (2) and (4) 2L -1 = -2L + 13 => 4L = 12 => L = 3 i.e A = 5 and K = 4 Kenny spent 4 hours, Andrea spent 5 hours and Lucas spent 3 hours

OpenStudy (anonymous):

so i write that whole answer?

OpenStudy (anonymous):

my solution is complete and simple, u just need the equations

OpenStudy (anonymous):

by the way, analyze the question. I think it's a little tricky. How many MORE hours did Lucas study? hmmm

OpenStudy (anonymous):

so i should just type out your answer?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can you look at my other questions too?